OpenStudy (anonymous):

Given the equation −2x − 13 = 8x + 7, which order of operations completely solves for x? Add 2x, then subtract 8x, lastly subtract 7 Add 2x, then add 13, lastly divide by 10 Subtract 8x, then add 13, lastly divide by −10 Subtract 8x, then add 2, lastly add 13

OpenStudy (anonymous):

C the third one, subtract 8x, then add 13, lastly divide by 10 -2x -13 = 8x +7 take away 8x -10x -13 = 7 then add 13 -10x = 20 divide by -10 x = -2